### v1

@@ 5.1 System representation
The investigated CSTR is taken from Klatt and Engell (1998). We consider the following reaction scheme
@@
$k_{1}$ $=\mu_{\beta}k_{0,\mathrm{ab}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{ab}}}{T_{\mathrm{R}}+273.15}\right),$ (61a)
$k_{2}$ $=k_{0,\mathrm{bc}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{bc}}}{T_{\mathrm{R}}+273.15}\right),$ (61b)
$k_{3}$ $=k_{0,\mathrm{ad}}\exp\left(\frac{-\mu_{\alpha}E_{\mathrm{A},\mathrm{ad}}}{T_{\mathrm{R}}+273.15}\right).$ (61c)
The parameters are defined as the nominal values in Klatt and Engell (1998). The coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ will be uncertain in the following investigations.
@@ 5.2 Agent-environment interaction
$R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})=$ $-10^{-2}\cdot\left(\frac{F_{\mathrm{prev}}-F}{F_{\mathrm{max}}-F_{\mathrm{min}}}\right)^{2}$
$-10^{-2}\cdot\left(\frac{\dot{Q}_{\mathrm{prev}}-\dot{Q}}{\dot{Q}_{\mathrm{max}}-\dot{Q}_{\mathrm{min}}}\right)^{2}$ (63c)
$R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})=$ $-100\max\left\{0,\boldsymbol{s}_{\mathrm{lb}}-\boldsymbol{s},\boldsymbol{s}-\boldsymbol{s}_{\mathrm{ub}}\right\}$ (63d)
where $R_{\mathrm{ref}}(\boldsymbol{s},\boldsymbol{a})$ encodes the tracking of the references, $R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})$ encodes the low variations in the actions, and $R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})$ encodes the constraints.
@@
The initial states $\boldsymbol{s}_{0}$ of the environment from which all episodes start are sampled uniformly from the feasible state space, so $\boldsymbol{s}_{0}\sim\mathcal{U}(\boldsymbol{s}_{\mathrm{lb}},\boldsymbol{s}_{\mathrm{ub}})$. Further, it is assumed that the coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ are uncertain. Both coefficients are sampled from the uniform distributions $\mu_{\alpha}\sim\mathcal{U}(0.95,1.05)$ and  $\mu_{\beta}\sim\mathcal{U}(0.95,1.05)$ at the beginning of each episode. The sampled realizations of $\mu_{\alpha}$ and $\mu_{\beta}$ are unknown to the agent, which means that the agent needs to learn a control policy that maximizes the cumulative reward but remains robust simultaneously.
The considered agent is a nominal MPC agent that uses the same model as described in (60) and (61), respects the constraints from Table 1, takes the negative reward from (63a) as stage cost, and predicts $N=20$ steps into the future. It is assumed that the parameters are known except for $\mu_{\alpha}$ and $\mu_{\beta}$ as they are uncertain. To improve the MPC policy, we choose to parameterize those coefficients as $\tilde{\mu}_{\alpha}$ and $\tilde{\mu}_{\beta}$ such that
$\tilde{\mu}_{\alpha}=(1+\theta_{\alpha})\cdot 0.95,\quad\tilde{\mu}_{\beta}=(1+\theta_{\beta})\cdot 1.$ (64)
The agent then needs to learn the parameters $\boldsymbol{\theta}^{\top}=(\theta_{\alpha},\theta_{\beta})$ such that the agent performs robust and not too conservative.
We approximate the action-value function using a NN. The following four steps are repeated for each update step in order to ensure a highly accurate Q-function approximation.
  1. 1.
    Sampling of $n_{\mathrm{IC}}$ initial states $\boldsymbol{s}_{0}$ and uncertain coefficient combinations $\mu_{\alpha},\mu_{\beta}$.
  2. 2.
    MPC closed-loop operation with current RL parameters $\boldsymbol{\theta}_{k}$ without exploration for all $n_{\mathrm{IC}}$ initial states, to approximate the state-value function $V_{\boldsymbol{\phi}}(\boldsymbol{s})\approx V^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})$.
  3. 3.
    Explore for all transitioned states $\boldsymbol{s}$ with exploratory action $\boldsymbol{a}_{\mathrm{exp}}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})+\mathcal{N}(0,\boldsymbol{\Sigma}_{\mathrm{exp}})$ and based on this approximate $Q_{\boldsymbol{\psi}}(\boldsymbol{s},\boldsymbol{a})\approx Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ locally around the current policy using the state-value function $V_{\boldsymbol{\phi}}(\boldsymbol{s})$.
  4. 4.
    Compute and apply the update according to the desired update rule.
All main training routines are based on the python toolbox do-mpc (Fiedler et al., 2023), which uses a CasADi (Andersson et al., 2019) backend and relies on Ipopt (Wächter and Biegler, 2006).
5.3 Convergence behavior and data-efficiency
We first investigate how our proposed Gauss-Newton approximation (21b) performs compared to the more elaborate Hessian approximation (20) and how it performs compared to the state-of-the-art first-order approach Adam (Kingma and Ba, 2015) with respect to convergence speed and data-efficiency. For all investigated scenarios below, both second-order approaches share the same learning rate of $\alpha=0.1$, gradient momentum of $\beta=0.75$, Hessian momentum of $\eta=0.9$, and initial eigenvalue estimate of $\omega^{-1}=10^{2}$. For the first-order optimizer Adam two different learning rates of $\alpha=10^{-3}$ and $\alpha=10^{-2}$ are investigated. The default Adam hyperparameters of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ are used for the first and second-order moments. Please refer to Kingma and Ba (2015) for more details.
The observed parameter trajectories during training for the initial parameter combination of $\theta_{\alpha,0}=\theta_{\beta,0}=0$ are shown in Figure 5.
Figure 5: Trajectories of $\theta_{\alpha}$ (x-axis) and $\theta_{\beta}$ (y-axis) throughout the policy optimization. All trajectories start at the same initial guess of $\theta_{\alpha,0}=\theta_{\beta,0}=0$. The coloring accounts for $J(\theta_{\alpha},\theta_{\beta})$. The markers denote each considered algorithm. Approx. Newton refers to (20). $\alpha$ denotes the respective learning rates.
Both second-order approaches show stable and fast convergence whereas both Adam approaches struggle. While Adam with a sufficiently small learning rate of $\alpha=10^{-3}$ shows stable but slow training progress, Adam with a larger learning rate of $\alpha=10^{-2}$ appears to have large but unstable improvement throughout the training process. We account the improved convergence behavior of both second-order approaches to the fact that $J(\theta_{\alpha},\theta_{\beta})$ has a significant different curvature in the $\theta_{\alpha}$ and $\theta_{\beta}$-dimensions, which can also be observed from the scales of the $\theta_{\alpha}$ and $\theta_{\beta}$-axis. Contrary to Adam, both second-order approaches can exploit the approximate Hessians to precondition the gradients by their curvature estimate. Adam however must rely on its fixed learning rate, which is limited by the largest element within the gradient. Adam therefore makes rapid progress into the $\theta_{\alpha}$-dimension but struggles along the $\theta_{\beta}$-dimension as the learning rate is limited by $\theta_{\alpha}$. From this, we conclude that second-order approaches are of special interest for MPC as agent for RL since the influence of introduced parameters is difficult to assess beforehand and may vary significantly for the different parameter configurations.
Table 2: Averaged performance metrics of the different training approaches over five initial guesses.
Approach Average time per iteration Number of iterations until convergence Total wall-clock time until convergence Averaged expected cumulative reward
after convergence
$\bar{t}\leavevmode\nobreak\ \mathrm{[min\,iter^{-1}]}$ $N_{\mathrm{conv}}\leavevmode\nobreak\ \mathrm{[iter]}$ $\bar{t}_{\mathrm{conv}}\leavevmode\nobreak\ \mathrm{[min]}$ $J(\theta_{\alpha},\theta_{\beta})$
Adam 1.45 >100 >145 -73.3
Approx. Newton 1.98 65 129 -62.5
Gauss-Newton (prop.) 1.85 65 120 -63.2
We further investigate the robustness of the proposed approaches with respect to the choice of initial guesses. Therefore, we run the training runs with the same specifications as above but for five randomly sampled initial guesses for $\theta_{\alpha,0}\sim\mathcal{U}(-0.05,0.07)$ and $\theta_{\beta,0}\sim\mathcal{U}(-0.75,0.1)$. Figure 6 shows the averaged expected cumulative reward (lines) and its standard deviation (shaded) for the investigated approaches for the five initial guesses. We omit the results for Adam with $\alpha=10^{-2}$ as it shows unsuccessful and unstable learning behavior. All three approaches start with the same averaged initial accumulated reward and robustly improve over the iterations. As observed in Figure 5, the training for Adam with $\alpha=10^{-3}$ is slower than for both second-order approaches. Even after 100 parameter updates, the performance is still not on the same level as both second-order approaches. Both second-order approaches on the other side make rapid improvement in the beginning. The Gauss-Newton approach appears to converge after roughly 65 iterations, whereas the approximate Newton approach still improves marginally.
Figure 6: Average expected cumulative reward (lines) and standard deviation (shaded) for five initial guesses of $\theta_{\alpha}$ and $\theta_{\beta}$ for the investigated policy optimization methods.
Finally we want to point out that not only the number of iterations can be reduced but also the total wall-clock time. During training we measure the computation time per iteration for each training run of the aforementioned agents. Table 2 lists the average wall-clock time per iteration, the number of iterations until convergence, the total wall-clock time until convergence, and the expected cumulative reward after convergence of the different approaches. Although the average computation time per iteration for the proposed Gauss-Newton approach is larger than for Adam, the Gauss-Newton approach only needs 65 iterations to converge to an average expected cumulative reward of $J(\theta_{\alpha},\theta_{\beta})=-63.2$, whereas Adam only achieves a performance of $J(\theta_{\alpha},\theta_{\beta})=-73.3$ after 100 iterations. This results in a total wall-clock time of 120 min for the Gauss-Newton approach and 145 min for Adam. Even in comparison to the approximate Newton approach, the Gauss-Newton approach obtains a 9 min lower total wall-clock time to achieve a similar performance. We expect these improvements to be even more pronounced for more complex systems and MPC policies with higher nonlinearity, which will be the subject of future research.
5.4 Influence of momentum $\beta$ and $\eta$
After we saw that the convergence behavior of the training process can be improved by the proposed Gauss-Newton approach to a similar performance as the considered approximate Newton approach, we also investigate the effect of the momentum on the training stability and training speed for different combinations of $\beta$ and $\eta$.
Figure 7: Expected cumulative reward for the Gauss-Newton approach for $\alpha=0.1$, $\omega^{-1}=10^{1}$ and differing $\beta$ and $\eta$
Figure 7 shows the evolution of the expected cumulative reward $J(\theta_{\alpha},\theta_{\beta})$ for the proposed Gauss-Newton approach with different combinations of $\beta$ and $\eta$. The learning rate and the initial eigenvalue estimate are kept at $\alpha=0.1$ and $\omega^{-1}=10^{1}$. We observe slow but stable convergence if $\eta$ is chosen large such as $\eta=0.99$ regardless of $\beta$ since the exponential moving average is dominated by old values and therefore also by its initialization matrix $\boldsymbol{D}_{0}$. The influence becomes less for smaller momentum such as $\eta=0.9$. Here, the updates are non-conservative from the very beginning and also remain stable. Even smaller or no momentum such as $\eta=0.5$ or $\eta=0$ can cause instability as it can be seen for $\beta=0.75$ and $\eta=0.5$, which diverges after 20 iterations, or $\beta=0.99$ and $\eta=0$, which caused an update that rendered the MPC’s NLP unsolvable. Therefore, the training was aborted. We conclude that stable training results are obtainable for various combinations of $\beta$ and $\eta$. Overall, the robustness of the training algorithm is insensitive with respect to the exact choice of $\eta$ as long as it is chosen sufficiently large. However, in order to also obtain fast convergence, $\eta$ should be chosen not too large.
Finally, we also study the stabilizing effect on the training of the momentum applied to the Hessian estimates at the example of $\eta=0.9$. Figure 8 shows the evolution of the estimated eigenvalues of the Hessian along the training process for the local estimates $\tilde{\boldsymbol{B}}_{k}$, the biased estimate $\boldsymbol{D}_{k}$, and the bias-corrected estimate $\hat{\boldsymbol{D}}_{k}$. It can be seen that the point estimates $\tilde{\boldsymbol{B}}_{k}$ are noisy and that the exponential moving average smoothens out this noise. This is highly important since especially the second eigenvalue varies over more than two orders of magnitude. Further, the bias correction accounts in both cases for the initialization bias enabling conservative regularized Hessian estimates (large eigenvalues) in the beginning. Along the training progress, $\boldsymbol{D}_{k}$ and $\hat{\boldsymbol{D}}_{k}$ converge to the same trajectory as derived in (44).
Figure 8: Eigenvalues of the local Hessian estimate $\tilde{\boldsymbol{B}}_{k}$, the biased averaged Hessian estimate $\boldsymbol{D}_{k}$, and bias-corrected averaged Hessian estimate $\hat{\boldsymbol{D}}_{k}$ for the Gauss-Newton approach over the training steps for $\eta=0.9$ and $\omega^{-1}=10$.

### v2

@@ 5.1 System representation
The investigated CSTR is taken from klattGainschedulingTrajectoryControl1998. We consider the following reaction scheme
@@
$k_{1}$ $=\mu_{\beta}k_{0,\mathrm{ab}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{ab}}}{R\,(T_{\mathrm{R}}+273.15)}\right),$ (64a)
$k_{2}$ $=k_{0,\mathrm{bc}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{bc}}}{R\,(T_{\mathrm{R}}+273.15)}\right),$ (64b)
$k_{3}$ $=k_{0,\mathrm{ad}}\exp\left(\frac{-\mu_{\alpha}E_{\mathrm{A},\mathrm{ad}}}{R\,(T_{\mathrm{R}}+273.15)}\right).$ (64c)
The parameters are defined as the nominal values in klattGainschedulingTrajectoryControl1998. The coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ will be uncertain in the following investigations.
@@ 5.2 Agent-environment interaction
$R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})=$ $-10^{2}\cdot\left(\frac{F_{\mathrm{prev}}-F}{F_{\mathrm{max}}-F_{\mathrm{min}}}\right)^{2}$
$-10^{2}\cdot\left(\frac{\dot{Q}_{\mathrm{prev}}-\dot{Q}}{\dot{Q}_{\mathrm{max}}-\dot{Q}_{\mathrm{min}}}\right)^{2}$ (66c)
$R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})=$ $-10^{2}\max\left\{0,\boldsymbol{s}_{\mathrm{lb}}-\boldsymbol{s},\boldsymbol{s}-\boldsymbol{s}_{\mathrm{ub}}\right\}$ (66d)
where $R_{\mathrm{ref}}(\boldsymbol{s},\boldsymbol{a})$ encodes the tracking of the references, $R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})$ encodes the low variations in the actions, and $R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})$ encodes the constraints as soft constraints. Note that a soft constraint formulation does not fully prohibit constraint violations. Slight constraint violations may still be possible even for the optimal policy.
@@
The initial states $\boldsymbol{s}_{0}$ of the environment from which all episodes start are sampled uniformly from the feasible state space, so $\boldsymbol{s}_{0}\sim\mathcal{U}(\boldsymbol{s}_{\mathrm{lb}},\boldsymbol{s}_{\mathrm{ub}})$. Further, it is assumed that the coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ in (64c) and (64a) are uncertain. Both coefficients are sampled from the uniform distributions $\mu_{\alpha}\sim\mathcal{U}(0.95,1.05)$ and $\mu_{\beta}\sim\mathcal{U}(0.95,1.05)$ for each transition. Additionally, the subsequent physical state $\boldsymbol{s}^{+}_{\mathrm{phys}}$ is disturbed with additive process noise $\boldsymbol{w}\sim\mathcal{N}(\cdot|\mu=0,\boldsymbol{\Sigma}=\mathrm{diag}(0.025,0.025,0.5,0.5))$. The sampled realizations of $\mu_{\alpha}$ and $\mu_{\beta}$ as well as the realization of the additive uncertainty $\boldsymbol{w}$ are unknown to the agent, which means that the agent needs to learn a control policy that maximizes the cumulative reward but remains robust simultaneously.
The considered agent is a nominal MPC agent that uses the same model as described in (63) and (64), respects the constraints from Table 1, takes the negative reward from (66a) as stage cost, and predicts $N=20$ steps into the future. The MPC parameterization will vary in the subsequent subsections. We want to emphasize that the MPC policy does not necessarily need to use the exact system model to predict the future. Practical alternatives such as data-driven surrogate models or linear systems can also be used and parameterized.
In order to investigate the effect of the proposed adaptive trust region constrained Gauss-Newton approach, we avoid to approximate the Q-function with the help of function approximators as these can significantly affect computation time and introduce other undesired side effects. We therefore implement the system, the reward and the MPC in a differentiable fashion such that the gradient and Hessian of the Q-function can be computed. The proposed approach is not constrained to differentiable models but can also be applied to the fully model-free setting. As with conventional first order approaches, an accurate approximation of the Q-function, optimally using a compatible function approximator as derived in works like silverDeterministicPolicyGradient2014 and kordabadQuasiNewtonCompatibleActorCritic2025, should be used if possible to avoid any systematic errors. All main training routines are based on the python toolbox do-mpc (fiedlerDompcFAIRNonlinear2023), which uses a CasADi (anderssonCasADiSoftwareFramework2019) backend and relies on Ipopt (wachterImplementationInteriorpointFilter2006).
5.3 Scalability with the number of parameters
We first investigate the scalability of the proposed adaptive trust region constrained Gauss-Newton approach with respect to the number of trainable parameters. We train three increasingly more parameterized MPC agents, ranging from a low parameterized agent with two parameters, over a medium parameterized agent with 13 parameters, up to a highly parameterized agent with 33 parameters. The low parameterization reflects a high level of certainty in almost all system parameters, whereas a high parameterization reflects as low level of certainty, allowing the RL algorithm to tune the parameters to come up with an optimized policy. The low parameterized agent considers the uncertain parameters $E_{A,\mathrm{ad}}$ and $k_{0,\mathrm{ab}}$ to be trainable, whereas the medium parameterized agent also considers other kinetic and thermodynamic parameters, and the setpoints within the objective to be trainable. The highly parameterized agent also parametrizes the terminal cost. Each introduced parameterization shares a similar structure of
$\tilde{\boldsymbol{p}}_{i,\boldsymbol{\theta}}=(1+\boldsymbol{\theta}_{i})c_{i}\boldsymbol{p}_{i}$ (67)
where $\tilde{\boldsymbol{p}}_{i}$ denotes the $i$-th MPC parameter that is modified, $\boldsymbol{p}_{i}$ denotes the $i$-th nominal MPC parameter, $c_{i}$ denotes a coefficient that models model error, and $\boldsymbol{\theta}_{i}$ is the $i$-th trainable policy parameter. The terminal cost $V_{\mathrm{f},\boldsymbol{\theta}}$ is learned according to the following parameterization
$V_{\mathrm{f},\boldsymbol{\theta}}=\left(\boldsymbol{x}_{N}-\tilde{\boldsymbol{x}}_{\mathrm{ref,f},\boldsymbol{\theta}}\right)^{\top}\tilde{\boldsymbol{P}}_{\boldsymbol{\theta}}^{\top}\tilde{\boldsymbol{P}}_{\boldsymbol{\theta}}\left(\boldsymbol{x}_{N}-\tilde{\boldsymbol{x}}_{\mathrm{ref,f},\boldsymbol{\theta}}\right)$ (68)
with cost matrix $\tilde{\boldsymbol{P}}_{\boldsymbol{\theta}}=\mathrm{diag}\left(0,0,10^{-2},10^{-2}\right)+\boldsymbol{P}_{\boldsymbol{\theta}}$, where $\boldsymbol{P}_{\boldsymbol{\theta}}\in\mathbb{R}^{4\times 4}$ is a matrix with RL parameters in each cell, and setpoint $\boldsymbol{x}_{\mathrm{ref,f}}^{\top}=(0.5,0.5,126,120)$. Detailed information on the parameterization levels is provided in Table 2.
Table 2: Parameters for the different parameterization levels.
Level $i$ $\boldsymbol{p}_{i}$ $c_{i}$ $i$ $\boldsymbol{p}_{i}$ $c_{i}$
Low 11 $E_{A,\mathrm{ad}}$ $0.95$ 22 $k_{0,\mathrm{ab}}$ 11
Med 33 $k_{0,\mathrm{bc}}$ 11 44 ${k}_{0,\mathrm{ad}}$ 11
55 $E_{A,\mathrm{ab}}$ $0.95$ 66 $E_{A,\mathrm{bc}}$ $1.05$
77 $H_{\mathrm{R,ab}}$ $1.05$ 88 $H_{\mathrm{R,bc}}$ $1.05$
99 $H_{\mathrm{R,ad}}$ $0.95$ $10$ $c_{p}$ $0.98$
$11$ $c_{p,\mathrm{K}}$ $1.02$ $12$ $T_{\mathrm{R,ref}}$ $1.0$
$13$ $T_{\mathrm{K,ref}}$ 11 - - -
High $14\ldots 29$ $\boldsymbol{P}_{\boldsymbol{\theta}}$ 11 $30\ldots 33$ $\boldsymbol{x}_{\mathrm{ref,f}}$ 11
Each agent is trained with the first order Adam optimizer, the adaptive trust region constrained second order approach with Hessian approximation according to (22), and the proposed adaptive trust region constrained Gauss-Newton approach. The Adam optimizer uses the default hyperparameters of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ with learning rate $\alpha=10^{-2}$. Larger learning rates lead to instabilities. Both second order approaches share the same hyperparameters of $\beta_{1}=0.75$, $\beta_{2}=0.999$, $\eta=0.9$ and $\alpha=0.1$. For each RL iteration $k$, the agents collect data from $N_{\mathrm{IC}}=200$ initial conditions with a fixed length of $N_{\mathrm{ep},i}=100$ steps.
Figure 5 shows the learning curves for the three differently parameterized MPC agents trained with the three approaches. All approaches show stable training. The proposed Gauss-Newton approach improves more rapidly than the Adam optimizer especially in the first iterations, showing that the Gauss-Newton approach can lead to an improved performance with less data. Simultaneously, the approximate Newton approach also shows stable improvements, however it shows mixed results regarding convergence speed. While it turns out to be faster than the Adam optimizer for the low and high parametrization, it improves slower in the scenario with medium parameterization. We account this to the fact that inaccurate intermediate Hessian estimates can adversely affect high quality updates. This deterioration then first needs to be recovered from, compromising the improved convergence speed.
Figure 5: Expected cumulative reward for increasingly more parameterized MPC schemes. The number in parentheses denotes the total number of trainable parameters.
Figure 6: Mean time per RL iteration for differently parameterized MPC schemes. The numbers above each bar show the mean value. The error bars denote three standard deviations. The number in parentheses denotes the total number of trainable parameters.
Figure 6 shows the average time per RL iteration for the investigated scenarios. For all three parameterizations, the Adam optimizer has the smallest iteration time, followed by the proposed Gauss-Newton approach, which takes about double the iteration time in the investigated scenario. In both approaches, the iteration time increases moderately with an increasing number of parameters. This stays in contrast to the approximate Newton approach, for which the iteration time increases drastically. We account the drastic increase of the iteration time to the fact that in order to compute the Hessian approximation according to (22), the second order NLP sensitivities need to be computed, which scales quadratically with the number of parameters. This is not the case for the proposed Gauss-Newton approximation.
We therefore conclude that the proposed Gauss-Newton approach offers a computationally cheap alternative to second order approaches with Hessian approximation that use second order NLP sensitivities such as (22) since the computation time scales better to higher parameterized policies. In addition to that, we see that the computation time of the Gauss-Newton approach scales similar to the computation time of the Adam optimizer. Although the iteration cost of the Gauss-Newton approach is higher in the presented case study, the number of iterations until a reasonable performance is achieved is less. This renders the proposed Gauss-Newton approach of special interest for settings in which the number of interactions between the agent and the environment are the limiting factor.
5.4 Influence of the parameter scaling
Next, we investigate how the curvature with respect to the parameters affects the learning stability. This is of high interest since it is typically difficult to assess the sensitivity of the objective with respect to the policy parameters and its rate of change beforehand. Moreover especially for MPC policies, parameters can have significantly different influences. This affects first order optimization approaches in particular, as the maximum learning rate is generally limited by the fastest changing element in the gradient vector.
For the following investigation, we adapt the MPC parameterization of $\tilde{E}_{A,\mathrm{ad}}$ by introducing the scaling factor $\xi$
$\tilde{E}_{A,\mathrm{ad}}=(1+\xi\,\theta_{1})\cdot 0.95\cdot E_{A,\mathrm{ad}}$ (69)
with $\xi\in\{0.1,1,10\}$ scaling the gradient and Hessian of the expected cumulative reward with respect to $\theta_{1}$. Choosing $\xi=0.1$ represents the scenario of a well scaled optimization problem, whereas $\xi=10$ represents a poorly scaled one. In order to assess the robustness of the proposed Gauss-Newton approach with respect to this parameter curvature, we ran training runs for five random initial guesses, sampled uniformly according to $\theta_{1,0}\sim\mathcal{U}(-0.1\xi^{-1},0.1\xi^{-1})$ and $\theta_{2,0}\sim\mathcal{U}(-0.5,0.5)$ for the low parameterization, and additionally $\theta_{3,0},\ldots,\theta_{13,0}\sim\mathcal{U}(-0.1,0.1)$ for the medium parameterized policy. Due to computational intractability of the approximate Newton method for the high parameterization, we omit the investigation for the highly parameterized policy. In order to guarantee comparability and fairness, all approaches within each parameterization level use the same initial guesses. All approaches use the same hyperparameters as in Section 5.3.
Table 3: Median (med.) and average (avg.) final performance (expected cumulative reward) of the agents, trained with Adam, Approx. Newton (AN) and Gauss-Newton (GN) for five different initial parameter guesses on the differently scaled low and medium policy parameterizations.
Adam AN GN
$\xi$ med. avg. med. avg. med. avg.
low $0.1$ 66.6-66.6 68.9-68.9 66.7-66.7 68.0-68.0 66.7-66.7 68.0-68.0
11 66.7-66.7 66.7-66.7 66.7-66.7 66.8-66.8 66.7-66.7 66.7-66.7
$10$ 99.5-99.5 86.6-86.6 66.7-66.7 66.7-66.7 66.7-66.7 66.7-66.7
medium $0.1$ 67.2-67.2 74.6-74.6 67.0-67.0 74.8-74.8 67.0-67.0 74.3-74.3
11 67.0-67.0 74.8-74.8 67.1-67.1 72.2-72.2 67.1-67.1 74.1-74.1
$10$ 71.6-71.6 71.6-71.6 67.0-67.0 67.3-67.3 67.0-67.0 67.0-67.0
Figure 7: Median expected cumulative reward (bold blue lines) and expected cumulative reward for five initial parameter guesses (gray lines) for the investigated policy optimization methods. Each row corresponds to a value of $\xi$. The left and right group of subplots represent the different parameterization levels (low and medium). Adam uses $\alpha=10^{-2}$. Both second order approaches use $\alpha=10^{-1}$.
Figure 7 shows the learning curves for each initial parameter guess (gray lines) and the median performance (bold blue lines) for three differently scaled low and medium parameterized agents. For the low parameterized policy, all three algorithms perform well $\xi=0.1$ and $\xi=1$. For $\xi=10$ both second order approaches converge robustly towards the optimal policy for all initial parameter guesses, whereas the Adam optimizer fails to find the optimal policy in three out of five cases. We found that choosing a smaller learning rate of $\alpha=10^{-3}$ remedied that behavior, however the learning progress was significantly slower compared to both second order approaches. The results for the medium parameterized policy, show a similar robust convergence for $\xi=0.1$ and $\xi=1$ for all but one initial guess. In early training stages, the median performance of both second order approaches improves more rapidly than Adam underlining the performance gains due to second order approaches. For $\xi=10$, both second order approaches converge robustly towards the optimal policy even for the challenging initial guess. On the other side, Adam also converges robustly, but to a different local optimum. The median and average final performance of the differently trained agents is summarized in Table 3. For all investigated scenarios the proposed Gauss-Newton approach achieves an average performance that is at least as good as Adam and even outperforms it most of the time, highlighting again the potential performance gains of second order approaches.
We conclude that both second order approaches can be more resilient with respect to the sensitivity of the objective with respect to the parameter scale. This can be of high interest especially when using MPC as policy in RL as the influence of the parameters on the closed loop performance can vary significantly depending on the parameter.
5.5 Comparison to deep reinforcement learning
Finally, we also compare the performance of the trained MPC agents with a neural network (NN) policy trained with the TD3 algorithm for deterministic policies (fujimotoAddressingFunctionApproximation2018). The implementation of the stable-baselines3 (SB3) toolbox is used (raffinStablebaselines3ReliableReinforcement2021). In order to guarantee a fair comparison, we performed a hyperparameter grid search. All agents are trained for $10^{6}$ time steps, which is equivalent to $10^{4}$ episodes. This is the same number of time steps as used for training of the agents in Section 5.3. The final choice of hyperparameters is given in Table 4.
Table 4: Final hyperparameters for the TD3 algorithm. All other hyperparameters use the SB3 default implementation.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
Activation gelu Batch size $1024$ Buffer size $10^{5}$
Policy $[64,64]$ Train frequency $200$ episodes Polyak coefficient $10^{-1}$
Q-function $[256,128]$ Gradient steps $500$ Policy delay $10$
Figure 8 shows the learning curve of the NN agent trained with the TD3 algorithm from SB3. The learning curve is put into reference to the learning curves of the low parameterized MPC agents trained as described in Section 5.3. The figure shows that the MPC agents are already well initialized. Still, the policy is fine-tuned during training, improving the performance even more. The NN policy on the other side are not that well initialized. Even after training, the final performance does not approach the performance of the initial MPC policies. The performance of the different policies is summarized in Table 5.
Figure 8: Expected cumulative reward over the observed training episodes for the considered MPC and NN agents.
Table 5: Expected cumulative reward after different numbers of episodes.
Episodes 0 5 0005\,000 10 00010\,000
Adam 91.49-91.49 67.38-67.38 66.99-66.99
Approx. Newton 91.49-91.49 66.85-66.85 66.64-66.64
Gauss-Newton 91.49-91.49 66.86-66.86 66.66-66.66
SB3 TD3 506.41-506.41 192.80-192.80 167.90-167.90
To this end, we also exemplary compare the closed loop trajectories obtained via the untrained MPC agent, the MPC agent trained via the Gauss-Newton approach and the NN agent trained with the TD3 algorithm. Figure 9 shows those closed loop trajectories. The untrained MPC agent transitions rapidly from the initial condition towards the temperature setpoints. However, it keeps a constant offset between the temperatures and the setpoints. The upper constraint for $C_{\mathrm{A}}$ is slightly violated. The trained MPC agent also transitions rapidly but, contrary to the untrained MPC agent, minimizes the offset between the temperatures and their setpoints. Further, the constraint violation is decreased. Still they are not fully eliminated as constraint violations are only penalized via soft constraints. The trained MPC agent manages to find a policy that is a trade-off-between constraint violations, fast and accurate setpoint tracking, and low oscillations in the actions. The trained MPC agent is therefore optimal with respect to the provided reward. Note that these constraint violations can be reduced if the penalty in (66d) is increased. Lastly, the NN policy transitions to the setpoints more conservatively than the MPC policies. The setpoints are tracked well and the constraint is not violated due to the conservative behavior, missing out on potential performance gains. In opposite to the MPC policies, the action trajectory is almost constant for $F_{\mathrm{in}}$ but slightly more oscillatory for $\dot{Q}$.
Figure 9: Closed-loop trajectories of the untrained MPC agent, the MPC agent trained via the proposed Gauss-Newton approach, and the TD3 agent.